# MobileRL: Advancing Mobile Use Agents with Adaptive Online Reinforcement Learning

### v1

@@ Abstract
Building general-purpose graphical user interface (GUI) agents has become increasingly promising with the progress in vision language models. However, developing effective mobile GUI agents with reinforcement learning (RL) remains challenging due to the heavy-tailed distribution of task difficulty and the inefficiency of large-scale environment sampling. We present an online agentic reinforcement learning framework MobileRL to enhance GUI agents in mobile environments. Its core component is the Difficulty-Adaptive GRPO (AdaGRPO) algorithm. In AdaGRPO, we design difficulty-adaptive positive replay and failure curriculum filtering to adapt the model to different task difficulties. We introduce the shortest-path reward adjustment strategy to reshape rewards concerning the task length in multi-turn agentic tasks. Those strategies jointly stabilize RL training, improve sample efficiency, and generate strong performance across diverse mobile apps and tasks. We apply MobileRL to two open models (Qwen2.5-VL-7B-Instruct and GLM-4.1V-9B-Base). The resultant MobileRL-9B model achieves state-of-the-art results in terms of success rates on both AndroidWorld (75.8%) and AndroidLab  (46.8%). The MobileRL framework is adopted in the AutoGLM products, and also open-sourced at https://github.com/THUDM/MobileRL.
@@
Figure 1: Task success rates on AndroidWorld (Rawles et al., 2024) and AndroidLab (Xu et al., 2024); hatched areas indicate gains from MobileRL on top of the SFT model.
@@ 1 Introduction
GUI agents—powered by vision language models—have enabled zero-shot interaction with web pages and mobile interfaces (Hong et al., 2023; OpenAI, ; Bai et al., 2025; Liu et al., 2024). To improve them, significant efforts have focused on supervised fine-tuning or offline imitation learning over static expert demonstrations (Rawles et al., 2023; Xu et al., 2024; Bai et al., 2024; Lu et al., 2025). However, these methods suffer from limited behavior coverage and poor error recovery (Chang et al., 2022).
Reinforcement learning (RL) with verifiable rewards presents a promising alternative (DeepSeek-AI et al., 2025; Hou et al., 2025). Yet, existing datasets with single-step expert labels  (Qin et al., 2025; Luo et al., 2025) are insufficient for training or evaluating policies on agentic tasks (i.e., planning and reasoning over multi-step action sequences). Although early progress has been made on online learning for GUI agents (Bai et al., 2024; Dong et al., 2025; Dai et al., 2025), efficiently scaling agentic RL in interactive mobile simulators remains largely unexplored.
@@
To address these challenges, we present an adaptive online agentic RL framework MobileRL for advancing mobile GUI agents. MobileRL consists of three components: reasoning-free supervised fine-turning (SFT), reasoning SFT, and agentic RL. The two SFT stages provide a warm-up for RL. Specifically, reasoning SFT enhances the handling of long and compositional instructions, reduces costly on-policy trials in mobile simulators, and enables the broad use of open or human-labeled datasets without relying on proprietary models.
@@
To enable effective online agentic RL, we introduce Difficulty–Adaptive Group Relative Policy Optimization (AdaGRPO). Built upon group relative policy optimization (GRPO) (Shao et al., 2024), its core idea is to adapt optimization to instance difficulty and explicitly reward solution efficiency. AdaGRPO designs three key strategies: (i) Difficulty-Adaptive Positive Replay (AdaPR) maintains a curated buffer of challenging, high-quality trajectories and balances them with fresh on-policy samples. In sparse-reward mobile environments, difficult successes are rare yet highly informative; replaying them amplifies their learning signal and stabilizes policy updates. (ii) Failure Curriculum Filtering (FCF) down-weights persistently unsolvable tasks using online difficulty statistics, reallocating computational budget toward challenging but feasible instances. Given the heavy-tailed difficulty distribution observed in mobile agent benchmarks (Xu et al., 2024; Rawles et al., 2024), pruning hard dead-ends improves sample efficiency while retaining signal from recoverable failures. (iii) Shortest-Path Reward Adjustment (SPA) reshapes the reward function based on completion length, assigning higher returns to shorter solutions. Length-sensitive rewards counteract bias toward verbose and better align with user preferences in mobile interaction contexts.
@@
We implement MobileRL in the AgentRL framework (Zhang et al., 2025), which supports multi-task, multi-turn agentic RL training. Unlike previous Android simulator implementations (Toyama et al., 2021; Rawles et al., 2024)—which generally do not support true concurrent execution, our framework sustains high throughput that orchestrates hundreds of Dockerized Android virtual devices (AVDs) across multiple machines. This setup enables concurrent interaction with over 1,000 environments while preserving reproducibility. Since most open-source benchmarks and simulators are built upon the Android operating system (Toyama et al., 2021; Rawles et al., 2024), this design ensures seamless compatibility and faithful reproduction of environment behaviors.
@@
We train MobileRL on Qwen2.5-VL-7B-Instruct (Bai et al., 2025) and GLM-4.1V-9B-Base (Team et al., 2025), producing MobileRL-7B and MobileRL-9B, respectively. First, MobileRL-9B lifts the success rates to 75.8% on AndroidWorld and 46.8% on AndroidLab, significantly outperforming previous state-of-the-art results (64.2% and 41.2%, respectively). Second, though MobileRL-7B is substantially smaller than the 72B-parameter UI-Tars-1.5 (Qin et al., 2025) and UI-Genie-Agent (Xiao et al., 2025), it achieves significantly better performance (e.g., +7.8% on AndroidWorld). Third, extensive ablation studies demonstrate the design of AdaGRPO. Finally, the MobileRL framework is also deployed in the AutoGLM (Liu et al., 2024) production pipeline.
@@ 2 MobileRL
We model the mobile GUI agent as a finite–horizon Markov Decision Process (MDP) (Littman, 2009), $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\mu_{0},H)$. The state space $\mathcal{S}$ comprises possible GUI states; each state $s_{t}$ contains a screenshot of the current device screen and its structured UI hierarchy extracted from XML metadata. The discrete action space $\mathcal{A}$ includes atomic operations such as Tap, Swipe, Type, LongPress, and Finish. The transition function $P(s_{t+1}\!\mid\!s_{t},a_{t})$ captures the dynamics induced by the Android OS and installed applications. Rewards are binary and terminal: the agent receives $1$ if the instruction is completed successfully and 0 otherwise.
Formally, at the start of an episode (t=0t{=}0) the agent receives an instruction $c\in\mathcal{C}$ and initial state $s_{0}\sim\mu_{0}$. At each timestep $t$, the agent observes $s_{t}\in\mathcal{S}$ and selects an action $a_{t}\in\mathcal{A}$ according to its conditional policy $\pi_{\theta}(a_{t}\mid s_{t},c)$. The environment then transitions to $s_{t+1}$. An episode yields a trajectory $\tau=(s_{0},a_{0},s_{1},\ldots,s_{T})$ and terminates upon successful completion of the instruction or when the horizon $H\in\mathbb{N}$ is reached ($T\leq H$). Because rewards are sparse and terminal, the return reduces to the final outcome $r_{T}=R(s_{T})\in\{0,1\}$, and the learning objective is: $\theta^{\star}=\arg\max_{\theta}\;\mathbb{E}_{(s_{0},c)\sim\mu_{0},\;\tau\sim\pi_{\theta}}\!\bigl{[}R(s_{T})\bigr{]}$.
To enable accurate action grounding and interaction with mobile interfaces, we define the agent’s observation space as a dual presentation: the current screenshot of the mobile phone and its corresponding compressed Extensible Markup Language (XML) information. The preprocessing steps for XML simplification are detailed in Appendix D.
@@ 2.1 The MobileRL Framework
In agentic RL training, sampling in virtual-device environments is usually inefficient; thus, starting online RL directly from a base model was found to be excessively time-consuming in preliminary experiments. Therefore, we perform SFT with expert demonstration data obtained by following the data collection protocol of (Xu et al., 2024). The curated instruction-tuning expert dataset covers 52 applications and 500k interaction steps. Note that this data is reasoning-free.
@@
To further construct a stronger reasoning policy initializer, we perform reasoning SFT via an iterative reasoning refinement strategy over the expert dataset. Manually collected expert demonstrations dataset for mobile use often contains only the final action sequence, omitting intermediate reasoning. Training solely on such “black-box” trajectories yields opaque policies, while many unlabeled tasks remain unused. We leverage an off-the-shelf instruction model to activate expert data and bootstrap a reasoning-augmented training set from raw demonstrations, yielding a structured and transparent policy initialization.
Concretely, we iteratively build the reasoning instruction–tuning pairs in three stages:
  • Stage 0 (Bootstrap sampling): For each task $x$ with expert answer $a^{*}$, the Instruct model $M$ generates diverse candidate reasoning–action pairs $(c_{k},a_{k})$ (via, e.g., temperature/nucleus sampling). Whenever $a_{k}=a^{*}$, we retain $(x,c_{k},a^{*})$ in $\mathcal{D}_{R}$.
  • Stage 1 (Supervised fine-tuning): Train an initial reasoning policy $\pi^{R}_{0}$ on $\mathcal{D}_{R}$.
  • Stage 2 (Iterative refinement): At iteration $t$, $\pi^{R}_{t}$ proposes candidates; those matching $a^{*}$ are scored by conciseness and $\log P(c)$. The best explanation $c^{*}$ is added to $\mathcal{D}_{\mathrm{new}}$, and $\pi^{R}_{t+1}$ is obtained by fine-tuning. We stop when the match rate saturates.
The resulting reasoning-oriented fine-tuning corpus (about 71.4k steps) is trained for two epochs to produce the reasoning warm-up model used for agentic RL training.
@@
Briefly, GRPO advances proximal policy optimization  (Schulman et al., 2017) by replacing the learned value baseline with an on-the-fly, group-relative baseline computed from a set of trajectories for the same task. Given a state $s$, we sample a group of $G$ actions $a_{1},\ldots,a_{G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid s)$ and define the group-relative advantage using the reward function $R$:
$A^{\mathrm{GRPO}}(s,a_{i})=\frac{R(s,a_{i})-\mathrm{avg}_{j=1}^{G}\,R(s,a_{j})}{\mathrm{std}_{j=1}^{G}\,R(s,a_{j})}\,.$
Let $T_{i}$ denote the number of steps for sample $i$ and $s_{i,t},a_{i,t}$ be the state and action at step $t$. Define the per-step importance ratio and the per-step advantage as
$\rho_{i,t}(\theta)=\frac{\pi_{\theta}(a_{i,t}\mid s_{i,t})}{\pi_{\theta_{\mathrm{old}}}(a_{i,t}\mid s_{i,t})},\qquad\hat{A}_{i,t}\;\equiv\;A^{\mathrm{GRPO}}(s_{i,t},a_{i,t}).$
The policy maximizes the clipped surrogate
$J^{\mathrm{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}s\sim d_{\pi_{\theta_{\mathrm{old}}}},\\ a_{1:G}\sim\pi_{\theta_{\mathrm{old}}}(\cdot\mid s)\end{subarray}}\Biggl{[}\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\min\!\Bigl{(}\rho_{i,t}(\theta)\,\hat{A}_{i,t},\;\mathrm{clip}\!\bigl{(}\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\bigr{)}\,\hat{A}_{i,t}\Bigr{)}\;$
$-\;\beta\,D_{\mathrm{KL}}\!\bigl{(}\pi_{\theta}\,\|\,\pi_{\theta_{\mathrm{ref}}}\bigr{)}\Biggr{]},$
where $D_{\mathrm{KL}}(p\|q)=\sum_{a}p(a)\log\frac{p(a)}{q(a)}$ and $d_{\pi_{\theta_{\mathrm{old}}}}$ denotes the state visitation distribution.
@@ 2.2 Difficulty–Adaptive GRPO
We develop Difficulty–Adaptive GRPO (AdaGRPO) with three strategies—shortest-path reward adjustment (SPA), difficulty–adaptive positive replay (AdaPR) and failure curriculum filtering (FCF)—to address the challenges faced in mobile agentic RL.
First, in multi-turn mobile agentic tasks—where immediate rewards are absent within a single round, unlike single-turn settings (Lu et al., 2025; Team et al., 2025)—the reward allocation strategy must be redesigned. Beyond assigning a uniform terminal reward, we introduce Shortest-Path Reward Adjustment (SPA), which reshapes rewards with respect to task length. The goal of the adjustment is to provide more informative learning signals, guiding the model toward accurate and efficient completion paths and facilitating the computation of trajectory-level advantages.
@@
2.2.1 Shortest-Path Reward Adjustment (SPA)

### v2

# MobileRL: Online Agentic Reinforcement Learning for Mobile GUI Agents

@@ Abstract
Building general-purpose graphical user interface (GUI) agents has become increasingly promising with the progress in vision language models. However, developing effective mobile GUI agents with reinforcement learning (RL) remains challenging due to the heavy-tailed distribution of task difficulty and the inefficiency of large-scale environment sampling. We present an online agentic reinforcement learning framework MobileRL to enhance GUI agents in mobile environments. Its core component is the Difficulty-ADAptive GRPO (AdaGRPO) algorithm. In AdaGRPO, we design difficulty-adaptive positive replay and failure curriculum filtering to adapt the model to different task difficulties. We introduce the shortest-path reward adjustment strategy to reshape rewards concerning the task length in multi-turn agentic tasks. Those strategies jointly stabilize RL training, improve sample efficiency, and generate strong performance across diverse mobile apps and tasks. We apply MobileRL to two open models (Qwen2.5-VL-7B-Instruct and GLM-4.1V-9B-Base). The resultant MobileRL-9B model achieves state-of-the-art results in terms of success rates on both AndroidWorld (80.2%) and AndroidLab (53.6%). The MobileRL framework is open-sourced at https://github.com/THUDM/MobileRL.
@@
Figure 1: Left and Center: Task success rates on AndroidWorld (Rawles et al., 2024) and AndroidLab (Xu et al., 2024); hatched areas indicate gains from MobileRL on top of the SFT model. Right: Trajectory-level success rate curves on AndroidWorld train and test sets during RL training.
@@ 1 Introduction
GUI agents—powered by vision language models—have enabled zero-shot interaction with web pages and mobile interfaces (Hong et al., 2023; OpenAI, ; Bai et al., 2025; Liu et al., 2024). To improve them, significant efforts have focused on supervised fine-tuning or offline imitation learning over static expert demonstrations (Rawles et al., 2023; Xu et al., 2024; Bai et al., 2024; Lu et al., 2025). However, these methods suffer from limited behavior coverage and poor error recovery (Chang et al., 2022). Reinforcement learning (RL) with verifiable rewards presents a promising alternative (DeepSeek-AI et al., 2025; Hou et al., 2025). Yet, existing datasets with single-step expert labels  (Qin et al., 2025; Luo et al., 2025) are insufficient for training or evaluating policies on agentic tasks (i.e., planning and reasoning over multi-step action sequences). Although early progress has been made on online learning for GUI agents (Bai et al., 2024; Dong et al., 2025; Dai et al., 2025), efficiently scaling agentic RL in interactive mobile simulators remains largely unexplored.
@@
To address these challenges, we present an adaptive online agentic RL framework MobileRL for advancing mobile GUI agents. MobileRL consists of three components: reasoning-free supervised fine-tuning (SFT), reasoning SFT, and agentic RL. The two SFT stages provide a warm-up for RL. Specifically, reasoning SFT enhances the handling of long and compositional instructions, reduces costly on-policy trials in mobile simulators, and enables the broad use of open or human-labeled datasets without relying on proprietary models.
@@
To enable effective online agentic RL, we introduce Difficulty–ADAptive Group Relative Policy Optimization (AdaGRPO). Built upon group relative policy optimization (GRPO) (Shao et al., 2024), its core idea is to adapt optimization to instance difficulty and explicitly reward solution efficiency. AdaGRPO designs three key strategies: (i) Difficulty-Adaptive Positive Replay (AdaPR) maintains a curated buffer of challenging, high-quality trajectories and balances them with on-policy samples (Mnih et al., 2015; Zha et al., 2019). In sparse-reward mobile environments, difficult successes are rare yet highly informative; replaying them amplifies their learning signal and stabilizes policy updates. (ii) Failure Curriculum Filtering (FCF), as a simplified version of curriculum learning (Matiisen et al., 2019; Narvekar et al., 2020), down-weights persistently unsolvable tasks using online difficulty statistics, reallocating computational budget toward feasible instances. Given the heavy-tailed difficulty distribution observed in mobile agent benchmarks (Xu et al., 2024; Rawles et al., 2024), pruning hard dead-ends improves sample efficiency while retaining signal from recoverable failures. (iii) Shortest‐Path Reward Adjustment (SPA) reshapes the reward function based on completion length, assigning higher returns to shorter solutions. Length-sensitive rewards counteract bias toward verbose and better align with user preferences in mobile interaction contexts.
@@
We implement MobileRL in a Verl-based framework (Sheng et al., 2024), which supports multi-task, multi-turn agentic RL training. Unlike previous Android simulator implementations (Toyama et al., 2021; Rawles et al., 2024)—which generally do not support true concurrent execution, our framework sustains high throughput that orchestrates hundreds of Dockerized Android virtual devices (AVDs) across multiple machines. This setup enables concurrent interaction with over 1,000 environments while preserving reproducibility. Since most open-source benchmarks and simulators are built upon the Android operating system (Toyama et al., 2021; Rawles et al., 2024), this design ensures seamless compatibility and faithful reproduction of environment behaviors.
@@
We train MobileRL on Qwen2.5-VL-7B-Instruct (Bai et al., 2025) and GLM-4.1V-9B-Base (Team et al., 2025), producing MobileRL-7B and MobileRL-9B, respectively. MobileRL-9B lifts the success rates to 80.2%  on AndroidWorld and 53.6%  on AndroidLab, significantly outperforming previous state-of-the-art results (64.2% and 41.2%, respectively). MobileRL-7B outperforms the much larger 72B-parameter models UI-Tars-1.5 (Qin et al., 2025) and UI-Genie-Agent (Xiao et al., 2025) (e.g., +16% on AndroidWorld), despite being substantially smaller in scale. Also, extensive ablation studies demonstrate the effectiveness in the design of AdaGRPO.
In summary, our contributions are as follows:
  • MobileRL Framework & scalable sampling: We develop MobileRL with a two-stage warm-up followed by online agentic RL for mobile GUI agents. We further establish a distributed sampling implementation that coordinates hundreds of Dockerized Android virtual devices (AVDs), enabling reproducible large-scale training on Android benchmarks.
  • AdaGRPO Algorithm: We introduce Difficulty–Adaptive Group Relative Policy Optimization (AdaGRPO), which extends GRPO with (i) AdaPR for replaying challenging successful trajectories, (ii) FCF for down-weighting persistently unsolved tasks, and (iii) SPA for length-sensitive reward shaping, thereby accounting for instance difficulty and solution efficiency.
  • Empirical Results: Training on Qwen2.5-VL-7B-Instruct and GLM-4.1V-9B-Base yields MobileRL-7B and MobileRL-9B. MobileRL-9B reaches 80.2% on AndroidWorld and 53.6% on AndroidLab, surpassing previous bests (64.2% / 41.2%).
@@ 2 MobileRL
We model the mobile GUI agent as a finite-horizon Markov Decision Process (MDP) (Littman, 2009) $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,H,\mu_{0})$. The state space $\mathcal{S}$ contains all possible GUI states; a concrete state at time $t$ is $s_{t}\in\mathcal{S}$, which comprises a screenshot of the device screen together with the structured UI hierarchy parsed from XML metadata (Cf. Appendix C). The action space $\mathcal{A}$ is a finite set of atomic GUI operations; an action at time $t$ is $a_{t}\in\mathcal{A}$, including primitives such as Tap, Swipe, Type, Long Press, Launch, Home, Back, Wait, and a terminating action Finish. $P(s_{t+1}\mid s_{t},a_{t})$ represents the stochastic transition mechanism of the Android OS and installed applications, capturing the uncertainty over possible next states. We consider a finite horizon $H\in\mathbb{N}$. The initial condition is drawn from a joint distribution $\mu_{0}$ over instruction–state pairs, i.e., $(s_{0},c)\sim\mu_{0}$. Here, $c$ is the natural-language task instruction provided once at the beginning of the episode (t=0t{=}0). At each timestep $t=0,1,\dots$, the agent observes only the current state $s_{t}$ and samples an action according to a policy $\pi_{\theta}$, i.e., $a_{t}\sim\pi_{\theta}(\cdot\mid s_{t},c)$. The environment then transitions to $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$. An episode yields a trajectory $\tau=\big((s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T},a_{T})\big)$ and terminates either when the agent intentionally selects Finish in a success state or when the horizon is reached, with $T\leq H$. The reward is assigned only after task completion, such that $R(\tau)=r(s_{T},a_{T})$ with $r(s_{T},a_{T})\in\{0,1\}$ indicating success (1) or failure (0). Consequently, learning maximizes the success probability: $\theta^{\star}=\arg\max_{\theta}\;\mathbb{E}_{(s_{0},c)\sim\mu_{0},\;\tau\sim\pi_{\theta}}\!\bigl[R(\tau)\bigr]$.
@@ 2.1 The MobileRL Framework
In agentic RL training, sampling in virtual-device environments is usually inefficient; thus, starting online RL directly from a base model was found to be excessively time-consuming. Therefore, we perform SFT with expert demonstration data obtained by following the data collection protocol of (Xu et al., 2024), supplemented with the training split of the publicly available AndroidControl dataset (Li et al., 2024). Note that this data is reasoning-free.
@@
To further construct a stronger reasoning policy initializer, we perform reasoning SFT via an iterative reasoning refinement strategy over the expert dataset. Manually collected expert demonstrations dataset for mobile use often contains only the final action sequence, omitting intermediate reasoning. Training solely on such “black-box” trajectories yields opaque policies. We leverage an off-the-shelf instruction model to bootstrap a reasoning-augmented training set from raw demonstrations, yielding a structured and transparent policy initialization. Concretely, we iteratively build the reasoning instruction–tuning pairs in three stages:
  • Bootstrap sampling: For each task $x$ with expert answer $a^{*}$, the Instruct model $M$ generates diverse candidate reasoning–action pairs $(c_{k},a_{k})$. Whenever $a_{k}=a^{*}$, we retain $(x,c_{k},a^{*})$ in $\mathcal{D}_{R}$.
  • Supervised fine-tuning: Train an initial reasoning policy $\pi^{R}_{0}$ on $\mathcal{D}_{R}$.
  • Iterative refinement: At iteration $t$, $\pi^{R}_{t}$ proposes candidates; those matching $a^{*}$ are scored by correctness. The best explanation $c^{*}$ is added to $\mathcal{D}_{\mathrm{new}}$, and $\pi^{R}_{t+1}$ is obtained by fine-tuning.
The resulting reasoning-oriented fine-tuning corpus is trained for two epochs to produce the reasoning warm-up model used for agentic RL training.
@@
Briefly, GRPO advances proximal policy optimization (Schulman et al., 2017) by replacing the learned value baseline with an on-the-fly, group-relative baseline computed from a set of trajectories for the same task. Given a shared initial condition $(s_{0},c)\sim\mu_{0}$, we sample a group of $G$ trajectories $\mathcal{G}=\{\tau_{1},\ldots,\tau_{G}\}$ by rolling out $\pi_{\theta_{\mathrm{sample}}}$. Let $T_{i}$ denote the number of steps for trajectory $\tau_{i}$, $s_{i,t},a_{i,t}$ be the state and action at step $t$, and the trajectory-level reward for $\tau_{i}$ be $R(\tau_{i})\in\{0,1\}$. We define the group-relative trajectory-level advantage for any step $(s_{i,t},a_{i,t})$ on trajectory $\tau_{i}$ as $\hat{A}_{i,t}\;\equiv\;\frac{R(\tau_{i})-\mu_{R}}{\sigma_{R}}\,$ , where $\mu_{R}$ and $\sigma_{R}$ are the mean and standard deviation of $\{R(\tau_{j})\}_{j=1}^{G}$. The GRPO loss can be written in empirical form as
$\mathcal{L}^{\mathrm{GRPO}}(\theta)=-\,\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\min\!\Bigl(\rho_{i,t}(\theta)\,\hat{A}_{i,t},\;\mathrm{clip}\!\bigl(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon\bigr)\,\hat{A}_{i,t}\Bigr)$ (1)
where $\rho_{i,t}(\theta)=\frac{\pi_{\theta}(a_{i,t}\mid s_{i,t})}{\pi_{\theta_{\mathrm{sample}}}(a_{i,t}\mid s_{i,t})}$ is the token-wise importance sampling (IS) ratio. We add the KL loss $\beta\,D_{\mathrm{KL}}\!\bigl(\pi_{\theta}\,\|\,\pi_{\theta_{\mathrm{ref}}}\bigr)$ to prevent the model from deviating too much from the prior distribution.
@@ 2.2 Difficulty–Adaptive GRPO
We develop Difficulty–Adaptive GRPO (AdaGRPO) with three strategies—shortest-path reward adjustment (SPA), difficulty–adaptive positive replay (AdaPR), and failure curriculum filtering (FCF)—to address the challenges faced in mobile agentic RL.
First, in multi‐turn mobile agentic tasks—where immediate rewards are absent within a single round, unlike single‐turn settings (Lu et al., 2025; Team et al., 2025)—the reward allocation strategy must be redesigned. Beyond assigning a uniform terminal reward, we introduce Shortest-Path Reward Adjustment (SPA), which reshapes rewards with respect to task length. The goal of the adjustment is to provide more informative learning signals, guiding the model toward accurate and efficient completion paths and facilitating the computation of trajectory-level advantages.
@@
2.2.1 Shortest‐Path Reward Adjustment (SPA)